OpenStudy (anonymous):

if cos theta = sqrt2/7 and 3pi/2 < theta < 2pi. what are the values of sin theta and tan theta

OpenStudy (anonymous):

\[if \cos \Theta = \sqrt{2}/2 and 3\pi /2 < \Theta < 2\pi\]

OpenStudy (anonymous):

well they give u cos right? cos is adjacent/hypotenuse so the adjacent side is 2 and the hypotenuse is 7. use the pythagorean theorm to get the sine and then use the sin/cos to get tan. the 3pi/2 thing is the boundaries where the answers can be

OpenStudy (anonymous):

i never did a problem like this so im lost

zepdrix (zepdrix):

\[\Large\rm \color{orangered}{\frac{3\pi}{2}\lt \theta\lt 2\pi}\]What quadrant does this put our angle \(\Large\rm \theta\) in crystelle? :)

OpenStudy (anonymous):

theta is just a angle where if u take the cos of that angle will give u sqrt 2/7

OpenStudy (anonymous):

u kinda need to know ur circle of radians to answer this question

OpenStudy (anonymous):

hm i think quad 1 since its all positive? i just learned this today so lol

OpenStudy (anonymous):

I think it is 4 quadrent

OpenStudy (anonymous):

google circle of quadrants and look for the 3pie/2 and 2pie

OpenStudy (anonymous):

ok !

OpenStudy (anonymous):

270 degrees

zepdrix (zepdrix):

4th quadrant? :) Ok great. So draw a triangle in the 4th quadrant. Use the information they gave you to label the sides.